OpenStudy (anonymous):

Line A has a slope of and passes through the point (3, 3). Line B has a slope of and passes through the point (-4, -2). At what point does line A intersect line B? Line A intersects line B at the point (, )

OpenStudy (anonymous):

for my sis

OpenStudy (anonymous):

Line A:3/2 Line B:-1/3

OpenStudy (haleyelizabeth2017):

If I am not mistaken (it's been a while since I did this), we need to find the slope intercept forms of these lines and then either substitute them or eliminate.

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

im 17 i completely forgot how to do this.......

OpenStudy (anonymous):

this is for my sister

OpenStudy (haleyelizabeth2017):

so just plug in the slope and points into y=mx+b and solve for b....(do this for both lines) then we can finish by substitution.

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

line A intersects line B at the point (-1, -3).

OpenStudy (haleyelizabeth2017):

Okay......

OpenStudy (anonymous):

i just did it on a piece of paper

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

thx tho

OpenStudy (haleyelizabeth2017):

Okay...sorry I couldn't be of anymore assistance.

OpenStudy (anonymous):

np

OpenStudy (triciaal):

do you need any more help?

OpenStudy (dtan5457):

What did you get for the intercept point? (your answer)?

OpenStudy (dtan5457):

Oh wait you got -1,-3, correct :)

OpenStudy (triciaal):

@dtan5457 i did not do the problem. Do you want to show what you did or wait for me to do it?

OpenStudy (dtan5457):

@triciaal the asker already got the answer themselves

OpenStudy (triciaal):

@dtan5457 so you were directing your question at the asker but did not specify. you responded to my question about needing more help. read the thread. ok.

OpenStudy (dtan5457):

@triciaal I responded to you because you responded to my question which I answered myself.

OpenStudy (mathstudent55):

This is the point-slope form of the equation of a line: \(y - y_1 = m(x - x_1)\) for a line with slope \(m\), and point \( (x_1, y_1) \) Use it twice, once for each line. Each time, substitute m, x1 and y1 with the slope and coordinates of the given point. Then you will get the equation of each line. The next step is to solve the two equations simultaneously as a system of equations. That will give you the point of intersection of the two lines.

OpenStudy (texaschic101):

yep ^^ thats how I would do it